OpenStudy (anonymous):

The area of a round flat parachute is 380 square feet. Find the radius of the parachute.

OpenStudy (anonymous):

ok so to find the area of a circle, we use the formula: \[pir\] ryou know this, right?

OpenStudy (anonymous):

sorry, that was supposed to say pie(r)^2

OpenStudy (anonymous):

\[\pi(r)^2\]

OpenStudy (anonymous):

yes i know that

OpenStudy (anonymous):

can you hold up for a sec

OpenStudy (anonymous):

ok, so this is like an average algebraic problem, except this time we need to solve for r(the radius) so we set up this equation like this: π(r)^2 =380

OpenStudy (anonymous):

π(r)^2 =380 r^2=380/π r^2= 120.95775674984 (calculator) \[r=\sqrt{120.957756}\] r=10.99 ( approximately)

OpenStudy (anonymous):

so the radius of the parachute is approximately 10.99